Each kg of dough takes 11 W h to develop into a gas holding structure. 50 % of this work input is dissipated as heat.

What is the temperature rise in the dough leaving the mixer if 830 kg of dough ingredients are fed into the mixer?

– Heat capacity of dough is 2.5 kJ /kg
Answer #1

Heat given to change temperature = m*S*deltaT

830*11*0.50*3600 = 830*2.5e3*deltaT

deltaT = [ 830*11*0.50*3600]/( 830*2.5e3)

= 7.92 K answer